# Activation-Informed Merging of Large Language Models

### v1

@@ 4.1 Selection of FT Expert LLMs
To understand how AIM reacts with different merging methods, we conduct experiments with merging different experts fine-tuned from the same base model. The set of experts we use is the same set of experts used by the two latest LLM merging algorithms in the literature, namely DARE Yu et al. (2024a) and WIDEN Yu et al. (2024b), which use the same three experts fine-tuned from Llama-2-13b Touvron et al. (2023). These experts include the WizardLM-13B Xu et al. (2024) model fine-tuned for instruction following, WizardMath-13B Luo et al. (2025) fine-tuned for superior mathematical reasoning, and llama-2-13b-code-alpaca which serves as the code expert Chaudhary (2023).
A Note On Weights: The weights we use in our experiments may not be exactly identical to the weights used in the experiments by Yu et al. (2024a) and Yu et al. (2024b), since the referenced weights for WizardMath-13B are no longer available publicly, instead we use a publicly available copy of the model. See Appendix A for more details.
@@ 4.3 Benchmarks Used For Evaluations
Given that the expert models we use in our experiments involve fine-tuning on instruction following, mathematical reasoning, and code generation we use several common benchmarks for each of these tasks. Specifically, we measure model performance on language understanding with the MMLU Hendrycks et al. (2021a) benchmark, instruction following with IFEval Zhou et al. (2023) benchmark, code generation with HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) benchmarks, and mathematical reasoning with the MATH Hendrycks et al. (2021b) and GSM8K Cobbe et al. (2021) benchmarks. For all benchmarks, we use the latest versions and up-to-date implementations developed by Gao et al. (2024) except for mathematical reasoning for which we use the chain of thought prompting used by Luo et al. (2025) to replicate the results of the original model as closely as possible. The code we use for these benchmark results will also be made publicly available for reproducibility.
In addition to the common benchmarks that we use to evaluate merged models, we also propose a new evaluation metric for LLM merging (or merging of different experts in general) which we believe helps better contextualize the value added by any given merging algorithm which we discuss in the following section.
@@ 4.4 Measuring Performance From an Optimization Perspective
We note that in most cases LLMs are not meant to operate as narrow expert models, unlike a large portion of deep learning applications where models are trained to perform very specific tasks such as classification or regression. LLMs in contrast, are generalist language models aiming to assist across a wide variety of tasks and applications. As such LLMs can be viewed from a multi-objective optimization perspective. In merging scenarios specifically, multiple expert models are brought together to create a merged model aiming to find the balance of performance across the different expertise of the fine-tuned models. In this sense, each expert can be thought of as optimized for a specific objective. This perspective lends itself rather well to a multi-objective optimization view of the problem. Given this, only looking at how each model performs on each of the benchmarks does not give us a full picture of the multi-objective goal of merging.
@@
Let $\mathbf{r}=(r_{1},r_{2},\dots,r_{N})$ denote the reference point in this space, which we set to $(0,0,\dots,0)$ to ensure hypervolume calculations are consistently defined. Given a set $S$ of FT models and the pre-trained base model, let $S^{*}\subseteq S$ denote the subset of models that are Pareto-optimal, i.e., models that are not dominated by any other model in $S$. The hypervolume of this set, denoted as $HV(S^{*})$, is defined as:
$HV(S^{*})=\lambda\left(\bigcup_{\mathbf{x}\in S^{*}}\text{dom}(\mathbf{x},\mathbf{r})\right),$ (7)
@@
When a merged model $M$ is introduced, the new set becomes $S^{\prime}=S\cup\{M\}$, and the updated Pareto-optimal set is denoted as $S^{\prime*}$. Given this set including the merged model we can measure the added value of the merged model from a multi-objective perspective as the normalized hypervolume gain (HV Gain) as a result of adding this merged model:
HV $\text{HV Gain}=\sqrt[d]{HV(S^{\prime*})-HV(S^{*})}$ (8)
@@ 5 Experiments & Results
In this section, we present our results on applying AIM to different merging methods as well as an ablation study on how the $\omega$ hyper-parameter affects the performance of each merging method.
@@ 5.1 AIM Applied to Various Merging Approaches
Figure 2: The Pareto fronts of models under different scenarios. Note that the points in these plots represent all models benchmarked in Table 1, however, to make the plots more readable, we only visualize the dominating points in each case. The measured increases in HV Gain when AIM is applied can be clearly seen in the Pareto frontier shifting further forward when AIM is applied compared to when only a population of merged models is evaluated.
@@
Table 1: Benchmark Results Across Various Merging Scenarios and Methods. Percentage changes are shown relative to models merged without AIM, with these differences highlighted for each metric. The highest-performing fine-tuned large language models and base models are highlighted in yellow, while the best-performing merged models are marked in blue. The results demonstrate that, in most cases, applying AIM significantly enhances the performance of merged models across all benchmarks, often increasing HV Gain.
@@ 5.2 Ablation study
To understand the effects of changing $\omega$ in AIM, we conduct an ablation study on the case of merging all three expert LLMs. For this study, we apply AIM with $\omega\in\{0.0,0.2,0.4,0.6,0.8\}$ and run the benchmarks on each merged model with each value of $\omega$. For brevity we do not report all benchmark results for each value here, instead, we track the hypervolume gain (The full set of results are presented in Appendix B). Specifically, to better visualize the effect of $\omega$ we measure the relative change in HV Gain compared to no AIM (i.e. $\omega=1.0$). We present these results in Figure 3. In most merging methods, we see that decreasing $\omega$ to even 00 still benefits the model performance while, in some cases, still increases the performance. However, in TIES merging particularly, we see that decreasing $\omega$ beyond $0.4$ seems to degrade performance, and setting $\omega$ to the most extreme case of $0.0$ does see some degradation in WIDEN as well. Given this, it seems that in these experiments, a value of $0.4$ balances the performance gains in methods responding well to AIM and the potential degradation of methods that benefit less from AIM. However, given this observation that in some cases pushing $\omega$ to 00 still yields benefits, there may be some value in exploring non-linear scaling of activation magnitudes and non-linear relaxation schemes that could further boost performance in some cases.
@@ 6 Conclusion and Outlook
Through extensive empirical evaluations across multiple merging methods and benchmark tasks, we demonstrated that AIM consistently improves performance, often yielding superior results in comparison to the original merging methods it was applied to. These results empirically confirm our hypothesis on the importance of the activation space. Notably, AIM boosted merged model performance by up to 40% in some cases, underscoring the crucial rule and the potential of activation information in merging methods. Our findings highlight the necessity of incorporating activation-informed strategies when merging multiple fine-tuned models.
@@ Appendix A Reproducibility Details
Aside from the checkpoint we provide our code and the link to the publicly available calibration data we use in our work. Our code is publicly available at https://github.com/ahnobari/ActivationInformedMerging.git and the calibration data can be found at https://huggingface.co/datasets/mit-han-lab/pile-val-backup.
Checkpoints:
We also provide the checkpoints for the main experiment on hugging face:

### v2

@@ 4.1 Selection of FT Expert LLMs
To understand how AIM reacts with different merging methods, we conduct experiments with merging different experts fine-tuned from the same base model. The set of experts we use is the same set of experts used by the two latest LLM merging algorithms in the literature, namely DARE Yu et al. (2024a) and WIDEN Yu et al. (2024b), which use the same three experts fine-tuned from Llama-2-13b Touvron et al. (2023). These experts include the WizardLM-13B Xu et al. (2024) model fine-tuned for instruction following, WizardMath-13B Luo et al. (2025) fine-tuned for superior mathematical reasoning, and llama-2-13b-code-alpaca which serves as the code expert Chaudhary (2023). See Appendix A for more details.
@@ 4.3 Benchmarks Used For Evaluations
Given that the expert models we use in our experiments involve fine-tuning on instruction following, mathematical reasoning, and code generation we use several common benchmarks for each of these tasks. Specifically, we measure model performance on language understanding with the MMLU Hendrycks et al. (2021a) benchmark, instruction following with IFEval Zhou et al. (2023) benchmark, code generation with HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) benchmarks, and mathematical reasoning with the MATH Hendrycks et al. (2021b) and GSM8K Cobbe et al. (2021) benchmarks. For all benchmarks, we use the latest versions and up-to-date implementations developed by Gao et al. (2024) except for mathematical reasoning, for which we use the chain of thought prompting used by Luo et al. (2025) to replicate the results of the original model as closely as possible. The code we use for these benchmark results will also be made publicly available for reproducibility.
In addition to the common benchmarks that we use to evaluate merged models, we also propose a new evaluation metric for LLM merging (or merging of different experts in general), which we believe helps better contextualize the value added by any given merging algorithm and which we discuss in the following section.
@@ 4.4 Measuring Performance From an Optimization Perspective
We note that in most cases, LLMs are not meant to operate as narrow expert models, unlike a large portion of deep learning applications where models are trained to perform very specific tasks such as classification or regression. LLMs, in contrast, are generalist language models aiming to assist across a wide variety of tasks and applications. As such, LLMs can be viewed from a multi-objective optimization perspective. In merging scenarios specifically, multiple expert models are brought together to create a merged model aiming to find the balance of performance across the different expertise of the fine-tuned models. In this sense, each expert can be thought of as optimized for a specific objective. This perspective lends itself rather well to a multi-objective optimization view of the problem. Given this, only looking at how each model performs on each of the benchmarks does not give us a full picture of the multi-objective goal of merging.
@@
Let $\mathbf{r}=(r_{1},r_{2},\dots,r_{N})$ denote the reference point in this space, which we set to $(0,0,\dots,0)$ to ensure hypervolume calculations are consistently defined. Given a set $S$ of FT models and the pre-trained base model, let $S^{*}\subseteq S$ denote the subset of models that are Pareto-optimal, i.e., models that are not dominated by any other model in $S$. The hypervolume of this set, denoted as $\operatorname{HV}(S^{*})$, is defined as:
$\operatorname{HV}(S^{*})=\lambda\left(\bigcup_{\mathbf{x}\in S^{*}}\text{dom}(\mathbf{x},\mathbf{r})\right),$ (6)
@@
When a merged model $M$ is introduced, the new set becomes $S^{\prime}=S\cup\{M\}$, and the updated Pareto-optimal set is denoted as $S^{\prime*}$. Given this set including the merged model, we can measure the added value of the merged model from a multi-objective perspective as the normalized hypervolume gain (HV Gain) as a result of adding this merged model:
HV $\text{HV Gain}=\sqrt[d]{\operatorname{HV}(S^{\prime*})-\operatorname{HV}(S^{*})}$ (7)
@@ 5 Experiments & Results
In this section, we present our results on applying AIM to different merging methods as well as an ablation study on how the $\omega$ hyperparameter affects the performance of each merging method. All experiments are run using 4 H100 GPUs, and each set of benchmarks takes roughly 15 minutes to run for each checkpoint.
@@ 5.1 AIM Applied to Various Merging Approaches
Figure 2: The Pareto fronts of models under different scenarios. Note that the points in these plots represent all models benchmarked in Table 1, for better readability, we only visualize the dominating points in each case. The measured increases in HV Gain when AIM is applied can be clearly seen in the Pareto frontier shifting further forward when AIM is applied compared to when only a population of merged models is evaluated.
Table 1: Benchmark Results Across Various Merging Methods. Percentage changes are shown relative to models merged without AIM. The highest-performing fine-tuned and base models are highlighted in yellow, and the best-performing merged models are marked in blue. The results demonstrate that applying AIM significantly enhances the performance of merged models.
@@ 5.2 Ablation study
To understand the effects of changing $\omega$ in AIM, we conduct an ablation study on the case of merging all three expert LLMs. For this study, we apply AIM with $\omega\in\{0.0,0.2,0.4,0.6,0.8\}$ and run the benchmarks on each merged model with each value of $\omega$. For brevity, we do not report all benchmark results for each value here, instead, we track the hypervolume gain (The full set of results are presented in Appendix B). Specifically, to better visualize the effect of $\omega$, we measure the relative change in HV Gain compared to no AIM (i.e. $\omega=1.0$). We present these results in Figure 3. In most merging methods, we see that decreasing $\omega$ to even 00 still benefits the model performance while, in some cases, still increases the performance. However, in TIES merging particularly, we see that decreasing $\omega$ beyond $0.4$ seems to degrade performance, and setting $\omega$ to the most extreme case of $0.0$ does see some degradation in WIDEN as well. Given this, it seems that in these experiments, a value of $0.4$ balances the performance gains in methods responding well to AIM and the potential degradation of methods that benefit less from AIM. However, given this observation that in some cases pushing $\omega$ to 00 still yields benefits, there may be some value in exploring non-linear scaling of activation magnitudes and non-linear relaxation schemes that could further boost performance in some cases.
@@ 6 Conclusion and Outlook
Through extensive empirical evaluations across multiple merging methods and benchmark tasks, we demonstrated that AIM consistently improves performance, often yielding superior results in comparison to the original merging methods it was applied to. These results empirically confirm our hypothesis on the importance of the activation space. Notably, AIM boosted merged model performance by up to 40% in some cases, underscoring the crucial role and the potential of activation information in merging methods. Our findings highlight the necessity of incorporating activation-informed strategies when merging multiple fine-tuned models.
@@ Appendix A Reproducibility Details
A Note On Weights: The weights we use in our experiments may not be exactly identical to the weights used in the experiments by Yu et al. (2024a) and Yu et al. (2024b), since the referenced weights for WizardMath-13B are no longer available publicly, instead we use a publicly available copy of the model.
@@
Aside from the checkpoint, we provide our code and the link to the publicly available calibration data we use in our work. Our code is publicly available at https://anonymous.4open.science/r/ActivationInformedMerging and the calibration data can be found at https://huggingface.co/datasets/mit-han-lab/pile-val-backup.
@@ Appendix B Ablation Detailed Results
Appendix C Sensitivity-Based Formulation Comparison
In this section, we present the results of running post-merging relaxation using the gradient-based sensitivity-based formulation discussed in the main body. We run the relaxation scheme based on gradients with a $\omega=0.4$, which we use in AIM. Table 3 shows how performance changes across benchmarks with gradient-based relaxation. We observe that the resulting performance boost remains very close to activation-based relaxation, with HV gain largely unchanged, with both gradient-based and pure activation-based boost trading blows evenly (see Table 1). Noting this and given that activations alone do not come with significant memory and computational cost, activation-based relaxation, which does not require computing gradients, is a more memory-efficient and computationally inexpensive process.
Table 3: Model performance comparison across different benchmarks, after relaxation applied using the gradients of the models with respect to the calibration data.
Method Model(s) Relaxation HumanEval MBPP MMLU MATH GSM8K IFEval HV Gain
Base Models
- Base - 17.07 27.80 52.18 0.70 4.20 25.10 -
- Code - 17.07 31.60 52.91 6.00 24.10 26.25 -
- Instruction Tuned - 26.83 34.80 53.41 7.50 43.40 35.67 -
- Math - 15.24 27.60 51.89 13.10 59.10 21.58 -
Merged Models
DARE Average Code + Instruction Tuned No 26.83 34.40 53.53 8.40 45.80 33.42 0.27
Code + Math No 16.46 28.60 51.96 15.10 64.70 22.02 0.23
Instruction Tuned + Math No 5.49 19.00 51.08 9.80 54.30 32.35 0.18
DARE Ties Instruction Tuned + Math No 8.54 23.80 51.39 9.20 54.10 33.89 0.20
Code + Instruction Tuned No 30.49 35.20 53.40 8.60 46.20 33.28 0.28
Code + Math No 17.07 27.40 51.92 14.90 63.60 22.53 0.23
Instruction Tuned + Math No 4.27 20.20 51.50 10.00 54.20 31.31 0.18
Code + Math No 18.29 28.60 52.10 15.00 64.70 21.92 0.24
Ties Merging Code + Math No 15.85 26.80 51.86 14.30 62.60 21.63 0.20
Instruction Tuned + Math No 28.05 34.60 54.45 8.70 44.70 34.04 0.23
Code + Instruction Tuned No 16.46 23.60 52.70 2.70 5.40 24.48 0.00
WIDEN Instruction Tuned + Math No 24.39 30.40 54.20 14.60 66.00 30.82 0.30
Code + Math No 17.07 29.40 53.35 14.20 64.40 24.02 0.24
Code + Instruction Tuned No 26.22 35.60 54.90 8.30 45.00 30.42 0.27